A plane travelled 1320 miles to Houston and back. The trip there was with the wind. It took 10 hours. The trip back was into the wind. It took 15 hours. What is the speed of the plane in still air? What is the speed of the wind?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

Plane speed= x mph
wind speed= y mph
with wind speed= (x+y)
against wind speed (x-y)
d= 1320
1320 / (x+y) = 15
divide by 15 1
88 / (x+y) =
(x+y) = 88
d= = 1320 ............1
1320 /(x-y)= 10 against wind
divide by 10
132 /(x-y) =
x - 1 y = 132 .............2
add up (1) & (2)
2 x = 220
/ 2
x= 110 mph speed of Plane in still air
plug value of x in (1)
we get y= -22 mph speed of wind

Answer by stanbon(75887)

A plane travelled 1320 miles to Houston and back. The trip there was with the wind. It took 10 hours. The trip back was into the wind. It took 15 hours. What is the speed of the plane in still air? What is the speed of the wind?
--------
With wind DATA:
distance = 1320 miles ; time = 10 hrs ; rate = d/t = 1320/10 = 132 mph
-----
Against wind DATA:
distance = 1320 miles ; time = 15 hrs ; rate = d/t = 1320/15 = 88 mph
----
Equations:
With wind rate: p + w = 132
Against wind::: p - w = 88
-----
Subtract and solve for "w":
2w = 44
wind = 22 mph
========================
Cheers,
Stan H.
===================